# Boosting Data Analytics with Synthetic Volume Expansion

### 07 Jan 2024 v2

@@ 4.3 Feature Relevance for Tabular Data
where $R_{m}(\cdot)$ denotes the empirical risk, evaluated on an inference sample $\tilde{\bm{Z}}_{1}^{(m,d)}$ in Step 1 and $\tilde{\bm{Z}}_{1}^{(m)}$ with $m=\hat{m}$ in Step 3 of Syn-Test, $\tilde{f}$ and $\tilde{f}_{S^{c}}$ denote the estimated predictive function function with and without $S$, and $SE$ denotes the standard error.
To mitigate the bias in $T$ due to CatBoost, we propose a Monte Carlo debiasing technique. This method refines an estimated null distribution from one sub-inference sample by utilizing the corresponding test statistic values from the other sub-inference sample. Specifically, we consider test statistic values $(T(\tilde{\bm{Z}}_{1}^{(m,d)}))_{d=1}^{D}$, as per (6), based on synthetic samples $\tilde{\bm{Z}}_{1}^{(m,d)}$ in Step 1, derived from the null generative model to approximate the null distribution, and the corresponding $(T({\tilde{\bm{Z}}}_{2}^{(m,d)}))_{d=1}^{D}$ to evaluate the Type-I error in Step 2, with ${\tilde{\bm{Z}}}_{2}^{(m,d)}$ being synthetic samples contingent on ${\mathcal{S}_{2}}$. The empirical CDF of $(T({\tilde{\bm{Z}}}_{1}^{(m,d)}))_{d=1}^{D}$, denoted as $\tilde{F}_{0}$, renders an approximated null distribution. To rectify the bias in $\tilde{F}_{0}$, we deploy the empirical CDF of $(T(\tilde{\bm{Z}}_{1}^{(m,d)})-D^{-1}\sum_{j=1}^{D}T({\tilde{\bm{Z}}}_{2}^{(m,d)}))_{d=1}^{D}$, which centers at zero under $H_{0}$. A visual depiction can be found in Figure 5 (middle row).
In (6), we reject $H_{0}$ if $T$ manifests as large. To compute the test statistic values in Steps 1 and 3, we generate an additional synthetic sample of size $2N$ and split it evenly into two subsamples. Using the first subsample, we train a CatBoost model $\operatorname{E}(Y|\bm{X})=f(\bm{X})$ to forecast $Y$ employing all features $\bm{X}$, resulting in $\tilde{f}$. In parallel, we train another CatBoost model $\operatorname{E}(Y|\bm{X}_{S^{c}})=f(\bm{X}_{S^{c}})$ using features $\bm{X}_{S^{c}}$, yielding $\tilde{f}_{S^{c}}$. By employing the synthetic sample to compute full and null predictive models, yielding $\tilde{f}$ and $\tilde{f}_{S^{c}}$, we can mitigate the intrinsic bias and asymptotics highlighted in (Dai, Shen and Pan, 2022) stemming from a limited inference size. This behavior is demonstrated in Figure 5 (middle row).
To refine a generative model under the null hypothesis, researchers often employ permutation by replacing redundant predictor vectors $\bm{X}_{S}$ with irrelevant values (Dai, Shen and Pan, 2022). However, this approach may not preserve the correlation structures between $\bm{X}_{S}$ and $\bm{X}_{S^{c}}$. Addressing this issue, we introduce a novel method that maintains these correlation structures while ensuring the feature irrelevance of $\bm{X}_{S}$ on $Y$ given the rest of the features. Our procedure involves two steps:
@@
    We then generate synthetic data tuples $(Y,\bm{X}_{S},\bm{X}_{S^{c}})$ using this model. Then modify these tuples by replacing $\bm{X}_{S}$ with the predicted values $g(\bm{X}_{S^{c}})$, resulting in new tuples $(Y,g(\bm{X}_{S^{c}}),\bm{X}_{S^{c}})$. This process ensures compliance with a specific subclass under the risk invariance of $H_{0}$ by creating conditionally independent tuples, as described by Dai, Shen and Pan (2022).
@@
Finally, we designate an MC size of $D=1,000$ for estimating both the null distribution and the Type-I error in Step 1 of Syn-Test. The parameters are set as $\alpha=0.05$, $\varepsilon=0.01$, and the optimal $\hat{m}$ will be tuned based on the ratios $m/n\in{1,2,\ldots,20}$.
@@ 4.3.1 Real-Benchmark Examples
To contrast Syn-Test with its traditional counterpart, consider the significance test in (6). When the finite-sample null distribution of $T$ is unknown, the asymptotic distribution of the test statistic may require stringent assumptions (Dai, Shen and Pan, 2022). To circumvent this, we use synthetic samples generated from TDM to approximate the null distribution, as in (Liu, Shen and Shen, 2023). Contrary to that approach, we refrain from using data perturbation, thus eliminating the requirement to maintain the rank property for privacy protection.
@@
To facilitate knowledge transfer, we initially pre-train a TDM using a random subset of $13,209$ observations. For significance testing in (5), we adapt the one-split black-box test statistic in (Dai, Shen and Pan, 2022) with a training sample of size $6,605$ and an inference sample of size $826$. To perform the Syn-Test, we follow the splitting scheme illustrated in Figure 4. In detail, we divide the training sample equally into two subsets, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, for fine-tuning purposes. Additionally, we utilize the inference sample $\mathcal{S}_{3}$ for both model training and fine-tuning.
Figure 4: Illustration of Syn-Test in comparison with traditional feature significance test for black-box models. For the traditional approach, one needs a training split to train the black box model independently (Dai, Shen and Pan, 2022), which we split evenly into two subsamples for fine-tuning.
@@
Knowledge Transfer Across Distinct Distributions. The Adult dataset concerns adult income for $32,650$ males and 16,192 females based on the $1994$ census, including six numerical and eight nominal features, for example, age, work class, final weight, the number of years in education, marital status, working hours per week, and native country; see Kohavi et al. (1996) for more details. Our objective is to test if age, the number of years in education, and working hours per week are relevant to predicting if an adult female’s income surpasses the threshold of $50K$ per year.
As Figure 6 depicts, the pronounced differences across genders exist in the distributions of categories like income, age, marital status, occupation, and relationship. A pertinent question arises: can adult male data augment the synthetic data generation and subsequent female analysis?
To harness this knowledge transfer, we pre-train a TDM solely on the adult-male data with $32,650$ observations. For hypothesis testing in (5), with a focus on adult females, we utilize a training sample of $2,700$, divided evenly into $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. Additionally, we use an inference sample $\mathcal{S}_{3}$ consisting of $300$. Following the Syn-Test approach as described in Figure 4, we use $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ for fine-tuning the pre-trained TDM and while using $\mathcal{S}_{3}$ to avoid overfitting in model training and fine-tuning.
Figures 6 and 7 demonstrate that the synthetic female data produced by the diffusion model aligns more with the adult-female dataset than with the adult-male dataset. This alignment is evident by the Fréchet Inception Distance (FID), which measures the distributional differences between the generated and raw data vectors under the Gaussian assumption. The 1- and 2-Wasserstein distances between the empirical distributions of the two datasets provide further evidence (note that FID is 2-Wasserstein distance under Gaussian assumption). Notably, as shown in Table 3, the male-focused pre-trained TDM, once fine-tuned with a smaller female dataset, crafts synthetic female data with a diminished margin of error compared to models pre-trained on the adult-male data and adult-female data with $2700$ individuals. This result confirms that leveraging pre-trained adult-male data with a somewhat distinct distribution can enhance the TDM’s generation precision for females. This empirical validation emphasizes the imperative of refining a pre-trained model to maximize knowledge transfer and achieve unparalleled accuracy.
Figure 5 highlights three key observations. First, the top two figures demonstrate that $\hat{m}=6n$ consistently controls the Type-I error of the Syn-Test across both scenarios. Second, the middle two figures suggest successful debiasing through synthetic data generation. Lastly, the bottom two plots exhibit a pronounced rightward shift in the test statistic distribution when comparing synthetic data to those derived from raw data. This shift signifies an enhancement in the test’s power, directly attributable to the augmentation of the sample size.
Table 3: Comparison of FID-scores, 1- and 2-Wasserstein distances between the actual female sample and various synthetic samples. The labels “raw”, “pre-trained”, and “fine-tuned” represent the raw data, synthetic data generated by a pre-trained model, and synthetic data produced by a fine-tuned model. The evaluation sample size for “Female (raw)” is $1,350$ while each of “Male (raw)”, “Male (pre-trained)” and “Female (fine-tuned)” is $32,650$.
FID (Gaussian) 1-Wasserstein 2-Wasserstein
Female (raw) vs Male (raw) 1.971 1.968 2.125
Female (raw) vs Male (pre-trained) 2.051 1.967 2.127
Female (raw) vs Female (fine-tuned) 0.249 1.170 1.399
@@
Figure 6: Marginal distribution comparison for female (blue), male (purple), and synthetic female datasets (red). Categorical features use normalized bar plots, while numerical ones employ kernel density plots.
@@ 4.3.2 Simulation
In this section, we evaluate the effectiveness of Syn-Test in controlling Type-I errors through simulation studies. We use the model in (4) with a modification: an additional feature, $\bm{X}_{8}$, is included but does not contribute to the model. Here, $\bm{X}=(\bm{X}_{1},\ldots,\bm{X}_{7},\bm{X}_{8})$ is distributed uniformly on $[0,1]^{8}$ ($\text{Uniform}(0,1)^{8}$), and $\bm{\epsilon}$ follows a normal distribution $N(0,0.2^{2})$. We assess the relevance of the feature $\bm{X}_{8}$ in prediction, thereby examining the control capability of Type-I error by Syn-Test. We will use the one-split test statistic proposed by (Dai, Shen and Pan, 2022) in Syn-Test, although we don’t rely on their asymptotic theory for the test statistic.
@@ 5 Data Privacy
To generate synthetic data satisfying (7), we may employ techniques like Differentially Private Stochastic Gradient Descent (DP-SGD). DP-SGD injects a certain amount of Gaussian noise into the gradient updates during model training, ensuring that the model satisfies the desired privacy guarantees (Abadi et al., 2016). Differentially-private diffusion models (Ghalebikesabi et al., 2023) is an example of a point. This method is particularly effective in generating differentially private synthetic data while maintaining the utility of the data.
@@ 6 Discussion: Future of Data Science
This article unveils the Syn paradigm—a novel approach to data analytics using high-fidelity synthetic data derived from real-world insights. By addressing challenges in traditional data analytics, such as data scarcity and privacy issues, this paradigm underscores that high-fidelity synthetic data can amplify the precision and efficiency of data analytics with sample size augmentation, as evidenced by the case studies herein. However, it is crucial to acknowledge the generational effect inherent to the Syn framework. Consequently, a fusion of reality and simulation is essential for unlocking the potential of synthetic data, offering fresh perspectives for both scientific and engineering domains.
@@
A systematic evaluation of the Syn framework across diverse applications is essential for advancing data science.
@@ Appendix A Proofs
Let $g(m)=C_{\bm{\theta}}m^{-\alpha}+m\text{LTV}$. The assumption that $\text{LTV}>\text{R}(\widehat{\bm{\theta}}(\tilde{\bm{Z}}^{(m_{0})}))/m^{*}$ implies that $g(m^{*})-C_{\bm{\theta}}(m^{*})^{-\alpha}>\text{R}(\widehat{\bm{\theta}}(\tilde{\bm{Z}}^{(m_{0})}))$. Moreover, note that $\text{Gr}^{(m)}\geq\text{Gr}^{(m^{*})}$ if $m>m^{*}$. Hence, when $m>m^{*}$,
@@
Thus, by the definition of $m_{0}$, that $m_{0}\leq m^{*}<+\infty$. This completes the proof. ∎
@@
Concerning the power gain, note that $\tilde{\phi}_{m,\alpha}-\phi_{n,\alpha}=\tilde{\phi}_{m,\alpha}-\phi_{m,\alpha}+\Delta$. By definition and assumption, $|\tilde{\phi}_{m,\alpha}-\phi_{m,\alpha}|\leq\text{TV}(\tilde{\bm{Z}}^{(m)},\bm{Z}^{(m)})=m\text{TV}(\tilde{F},F)<\Delta$, indicating that $\tilde{\phi}_{m,\alpha}-\phi_{n,\alpha}>0$.
This completes the proof. ∎

### 10 Mar 2024 v3

@@ 3.3 Feature Relevance for Tabular Data
where $R_{m}(\cdot)$ denotes the empirical risk, evaluated on an inference sample $\tilde{\bm{Z}}_{1}^{(m,d)}$ in Step 1 and $\tilde{\bm{Z}}_{1}^{(m)}$ with $m=\hat{m}$ in Step 3 of Syn-Test, $\tilde{f}$ and $\tilde{f}_{S^{c}}$ denote the estimated predictive function function with and without $S$, and $SE$ denotes the standard error. Note that a large value of $T$ indicates the relevance of tested features to the response.
In (6), we reject $H_{0}$ if $T$ manifests as large. To compute the test statistic values in Steps 1 and 3, we generate an additional synthetic sample of size $2N$ and split it evenly into two subsamples. Using the first subsample, we train a CatBoost model $\operatorname{E}(Y|\bm{X})=f(\bm{X})$ to forecast $Y$ employing all features $\bm{X}$, resulting in $\tilde{f}$. In parallel, we train another CatBoost model $\operatorname{E}(Y|\bm{X}_{S^{c}})=f(\bm{X}_{S^{c}})$ using features $\bm{X}_{S^{c}}$, yielding $\tilde{f}_{S^{c}}$. By employing the synthetic sample to compute full and null predictive models, yielding $\tilde{f}$ and $\tilde{f}_{S^{c}}$, we can mitigate the intrinsic bias and asymptotics highlighted in (Dai, Shen and Pan, 2024) stemming from a limited inference size. This behavior is demonstrated in Figure 8 (middle row).
To refine a generative model under the null hypothesis, researchers often employ permutation by replacing redundant predictor vectors $\bm{X}_{S}$ with irrelevant values (Dai, Shen and Pan, 2024). However, this approach may not preserve the correlation structures between $\bm{X}_{S}$ and $\bm{X}_{S^{c}}$. Addressing this issue, we introduce a novel method that maintains these correlation structures while ensuring the feature irrelevance of $\bm{X}_{S}$ on $Y$ given the rest of the features. Our procedure involves two steps:
@@
    We generate synthetic data tuples $(Y,\bm{X}_{S},\bm{X}_{S^{c}})$ using this model. Then, we modify these tuples by replacing $\bm{X}_{S}$ with the predicted values $g(\bm{X}_{S^{c}})$, resulting in new tuples $(Y,g(\bm{X}_{S^{c}}),\bm{X}_{S^{c}})$. This process ensures compliance with a specific subclass under the risk invariance of $H_{0}$ by creating conditionally independent tuples, as described by Dai, Shen and Pan (2024).
@@
Finally, we designate an MC size of $D=1,000$ for estimating both the null distribution and the Type-I error in Step 1 of Syn-Test. The parameters are set as $\alpha=0.05$, $\varepsilon=0.01$, and the optimal $\hat{m}$ will be tuned based on the ratios $m/n\in\{1,2,\ldots,20\}$.
@@ 3.3.1 Real-Benchmark Examples
To contrast Syn-Test with its traditional counterpart, consider the significance test in (6). When the finite-sample null distribution of $T$ is unknown, the asymptotic distribution of the test statistic may require stringent assumptions (Dai, Shen and Pan, 2024). To circumvent this, we use synthetic samples generated from TDM to approximate the null distribution, as in (Liu, Shen and Shen, 2023). Contrary to that approach, we refrain from using data perturbation, thus eliminating the requirement to maintain the rank property for privacy protection.
@@
To facilitate knowledge transfer, we initially pre-train a TDM using the pre-training split with $13,209$ observations. For significance testing in (5), we adapt the one-split black-box test statistic in (Dai, Shen and Pan, 2024) with a training sample of size $6,605$ and an inference sample of size $826$. To perform the Syn-Test, we follow the splitting scheme illustrated in Figure 7. In detail, we divide the training sample equally into two subsets, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, for fine-tuning purposes. Additionally, we utilize $\mathcal{S}_{3}$ as a validation sample for both model training and fine-tuning. More details on the Syn-Test method can be found in Section 2.3.
Figure 7: Illustration of Syn-Test in comparison with traditional feature significance test for black-box models. For the traditional approach, one needs a training split to train the black box model independently (Dai, Shen and Pan, 2024), which we split evenly into two subsamples for fine-tuning. More details can be found in Section 2.3.
@@
Knowledge Transfer Across Distinct Distributions. In the study of the Adult dataset (Kohavi et al., 1996), we conduct a significance test to assess the relevance of three features—age, education years, and weekly working hours—in predicting whether an adult female’s income exceeds \$50K annually.
To enhance the test’s power for females, we generate synthetic data for females through knowledge transfer from the male dataset. This methodology aligns with the approach in the latter part of Section 3.2.1. Here, we pre-train a TDM on a dataset comprising $32,650$ adult male records. For our hypothesis testing in (5) focused on adult females, we use a training set of $2,700$, split equally between $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. Additionally, we use an inference sample of $300$ that also serves as a validation set $\mathcal{S}_{3}$ in the Syn-Test. This approach, as outlined in Section 2.3 and illustrated in Figure 7, involves fine-tuning the pre-trained TDM with $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ while utilizing $\mathcal{S}_{3}$ to prevent overfitting during model training and refinement. Table 3 and Figures 3 and 4 show the effectiveness of the fine-tuned TDM and address the importance of knowledge transfer for reducing generation error.
@@
Figure 8 (right column) highlights two key observations. First, the top figure demonstrates that $\hat{m}=6n$ consistently controls the Type-I error estimated from Syn-Test. Second, the bottom plot exhibits a pronounced rightward shift in the test statistic distribution when comparing synthetic data to those derived from raw data. This shift indicates an enhancement in the test’s power, directly attributable to the augmentation of the sample size.
@@ 3.3.2 Simulation
In this section, we evaluate the effectiveness of Syn-Test in controlling Type-I errors through simulation studies. We use the model in (4) with a modification: an additional feature, $\bm{X}_{8}$, is included but does not contribute to the model. Here, $\bm{X}=(\bm{X}_{1},\ldots,\bm{X}_{7},\bm{X}_{8})$ is distributed uniformly on $[0,1]^{8}$ ($\text{Uniform}(0,1)^{8}$), and $\bm{\epsilon}$ follows a normal distribution $N(0,0.2^{2})$. We assess the relevance of the feature $\bm{X}_{8}$ in prediction, thereby examining the control capability of Type-I error by Syn-Test. We will use the one-split test statistic proposed by (Dai, Shen and Pan, 2024) in Syn-Test, although we don’t rely on their asymptotic theory for the test statistic.
@@ 4 Data Privacy
To generate synthetic data satisfying (7), we may employ techniques like Differentially Private Stochastic Gradient Descent (DP-SGD). DP-SGD injects a certain amount of Gaussian noise into the gradient updates during model training, ensuring that the model satisfies the desired privacy guarantees (Abadi et al., 2016). Differentially-private diffusion models (Ghalebikesabi et al., 2023) is an example of a point. Further exploration is warranted.
@@ 5 Discussion: Future of Data Science
This article unveils the Syn paradigm—a novel approach to data analytics using high-fidelity synthetic data derived from real-world insights. By addressing challenges in traditional data analytics, such as data scarcity and privacy issues, this paradigm underscores that high-fidelity synthetic data can amplify the precision and efficiency of data analytics with sample size augmentation, as evidenced by the case studies herein. However, it is crucial to acknowledge the generational effect inherent to the Syn framework. Consequently, a fusion of raw and synthetic data is essential for unlocking their potential, offering fresh perspectives for both scientific and engineering domains.
In this study, although our primary focus is on tabular data—the predominant form in statistical analysis—we emphasize that the Syn framework’s utility extends across disparate types of data given available tools such as multimodal diffusion models (Xu et al., 2023). Its utility is reinforced by recent progress in generative modeling for vision, natural language, and multimodal data processing. This adaptability allows the Syn framework to perform comprehensive analyses of complex, multimodal data, such as electronic health records, which typically include a mix of standard tabular data, image scans, and unstructured text.
The importance of synthetic data in driving progress in data science and AI is increasingly recognized. However, a comprehensive evaluation of the Syn framework across a broad spectrum of applications is imperative. One key strategy for assessing the synthetic sample’s fidelity or reliability is comparing it with an independent validation sample that reflects the original data’s distribution. This assessment can be conducted through (i) exploratory analysis, like histograms and correlation matrices for visual inspection, and (ii) distributional distances, such as the Wasserstein distance, for a quantitative evaluation. Furthermore, in supervised learning tasks, the model’s performance can be scientifically validated using cross-validation metrics. Overall, validating generative models for more complex applications presents unique challenges, making the exploration of model robustness and data privacy critical areas for future research.
@@ Appendix A Proofs
Let $\text{LTV}\doteq f(\text{TV}(\tilde{F},F))$. Let $g(m)=C_{\bm{\theta}}m^{-\alpha}+m\text{LTV}$. The assumption that $\text{LTV}>\text{R}(\widehat{\bm{\theta}}(\tilde{\bm{Z}}^{(m_{0})}))/m^{*}$ implies that $g(m^{*})-C_{\bm{\theta}}(m^{*})^{-\alpha}>\text{R}(\widehat{\bm{\theta}}(\tilde{\bm{Z}}^{(m_{0})}))$. Moreover, note that $\text{Gr}^{(m)}\geq\text{Gr}^{(m^{*})}$ if $m>m^{*}$. Hence, when $m>m^{*}$,
@@
This completes the proof of (ii). Moreover, by the definition of $m_{0}$ and (ii), $m_{0}\leq m^{*}<+\infty$. This completes the proof of (i). ∎
@@
Concerning the power gain, note that $\tilde{\phi}_{m,\alpha}-\phi_{n,\alpha}=\tilde{\phi}_{m,\alpha}-\phi_{m,\alpha}+\Delta$. By definition and assumption, $|\tilde{\phi}_{m,\alpha}-\phi_{m,\alpha}|\leq\text{TV}(\tilde{\bm{Z}}^{(m)},\bm{Z}^{(m)})=m\text{TV}(\tilde{F},F)<\Delta$, indicating that $\tilde{\phi}_{m,\alpha}-\phi_{n,\alpha}>0$. This completes the proof. ∎
Appendix B Additional Experiment
In this section, we conduct a simulation study on traditional tabular data to demonstrate the efficacy of the Syn-Slm method, as discussed in Section 2.4. Utilizing the model in equation (4) as the basis for our data generation, we create training data of $1,000$ instances and test data of $500$ instances. For traditional analysis, we train CatBoost regression using the training dataset. For the Syn-Slm approach, a pre-trained TDM, initially pre-trained with pre-training data of size $2,000$, is fine-tuned on the training set. This fine-tuned model is then used for conditional generation (Zhang et al., 2023b) to approximate $\operatorname{E}(\bm{Y}|\bm{X})$ via a Monte Carlo method. Specifically, for any given input $\bm{x}$, we generate a series of $(\tilde{\bm{Y}}^{(d)})_{d=1}^{D}$ from the conditional distribution $\tilde{\bm{Y}}|\bm{X}=\bm{x}$ as learned by the TDM, and compute the estimate as $D^{-1}\sum_{d=1}^{D}\tilde{\bm{Y}}^{(d)}$. A notable advantage of Syn-Slm is its flexibility in handling various supervised learning tasks simultaneously without needing to retrain for different loss functions. For instance, one may use $\tilde{\bm{Y}}|\bm{X}=\bm{x}$ to estimate prediction intervals (Liu, Shen and Shen, 2023), quantiles, or other statistical measures with relative ease. The performance of both methods is assessed on the test dataset using RMSE as the metric. As indicated in Table 5, the Syn-Slm approach consistently outperforms the traditional CatBoost method across various noise levels $\sigma$, showcasing its superior predictive capability.
Table 5: RMSE of CatBoost and Syn-Slm across different noise magnitude $\sigma\in\{0.1,~{}0.2,~{}0.5,~{}1.0\}$. The number inside the parenthesis represents standard error from repeated conditional generations.
$\sigma=0.1$ $\sigma=0.2$ $\sigma=0.5$ $\sigma=1.0$
CatBoost 0.1227 0.2245 0.5271 1.0437
Syn-Slm 0.1202 (0.0005) 0.2171 (0.0008) 0.5148 (0.0020) 1.0161 (0.0038)
@@ Appendix B Additional Experiment
{supplement}
The code to reproduce the results presented in this paper is available at https://github.com/yifei-liu-stat/syn.
@@
[…]
Prior to input into the diffusion model, tabular data undergoes a transformation process. Continuous variables are normalized through quantile transformation, while categorical variables receive one-hot encoding. These processed variables are then concatenated to form the diffusion model’s initial input vector.
For training the models in our examples, we employ a FNN as the backbone of the diffusion model, along with a cosine scheduler (Nichol and Dhariwal, 2021) and the AdamW optimizer (Loshchilov and Hutter, 2017) for optimization. In real-world scenarios involving eight benchmark datasets in Sections 3.2.1 and 3.3.1, we develop pre-trained models using the configurations in Table 7. Model fine-tuning followed the same architectural design but typically involved lighter optimization — a reduced learning rate and fewer training iterations. For the simulation examples in Sections 3.2.1 and 3.3.1, we derive pre-trained models using the same pre-training configuration as that on the California dataset but with $50,000$ training iterations. Fine-tuning in these cases was carried out with $1,000$ iterations and a learning rate of $3\times 10^{-6}$.
Table 7: Model design and training details on eight real-world benchmark datasets. Adult-Male uses the same pre-training configuration as Adult does.
Dataset FNN dimensions Diffusion timestamps Iterations Learning rate Batch size
FB comments (512, 1024) 1000 30000 $6.3\times 10^{-4}$ 4096
Adult (256, 1024, 1024, 1024, 1024, 256) 100 30000 $2.0\times 10^{-3}$ 4096
House (128, 512, 512, 512, 512, 256) 1000 30000 $1.4\times 10^{-3}$ 4096
California (512, 256, 256, 256, 256, 128) 1000 30000 $1.3\times 10^{-4}$ 4096
Churn2 (512, 1024, 1024, 1024, 1024, 512) 100 30000 $8.1\times 10^{-4}$ 4096
Gesture (128, 512, 512, 1024) 1000 30000 $2.8\times 10^{-3}$ 4096
Abalone (512, 128) 100 30000 $3.4\times 10^{-4}$ 4096
Insurance (256, 512, 512, 512, 512, 256) 100 30000 $1.1\times 10^{-3}$ 4096